A satellite has a mass of 103 kg and is located at 2.01*10^6 m above the earth surface. what is the potential energy associated with the satellite at this location? (J) What is the magnitude of the gravitational force on the satellite? (N)
Answer #1

here,

mass of satellite , m1 = 103 kg

mass of earth , m2= 5.97 * 10^24 kg

the radius of earth , R = 6.371 * 10^6 m

the height above the surface , A = 2.01 * 10^6 m

the potential energy associated with the satellite at this location , PE = G * m1 * m2 / ( R + A)

PE = 6.67 * 10^-11 * 103 * 5.97 * 10^24 /(6.371 * 10^6 + 2.01 * 10^6)

PE = 4.89 * 10^9 J

the magnitude of the gravitational force on the satellite , F = G * m1 * m2 / ( R + A)^2

F = 6.67 * 10^-11 * 103 * 5.97 * 10^24 /(6.371 * 10^6 + 2.01 * 10^6)^2

F = 584 N